# Learning to Play Piano in the Real World

### v2

@@ Abstract
Towards the grand challenge of achieving human-level manipulation in robots, playing piano is a compelling testbed that requires strategic, precise, and flowing movements. Over the years, several works demonstrated hand-designed controllers on real world piano playing, while other works evaluated robot learning approaches on simulated piano scenarios. In this paper, we develop the first piano playing robotic system that makes use of learning approaches while also being deployed on a real world dexterous robot. Specifically, we make use of Sim2Real to train a policy in simulation using reinforcement learning before deploying the learned policy on a real world dexterous robot. In our experiments, we thoroughly evaluate the interplay between domain randomization and the accuracy of the dynamics model used in simulation. Moreover, we evaluate the robot’s performance across multiple songs with varying complexity to study the generalization of our learned policy. By providing a proof-of-concept of learning to play piano in the real world, we want to encourage the community to adopt piano playing as a compelling benchmark towards human-level manipulation. We open-source our code and show additional videos at www.lasr.org/research/learning-to-play-piano.
@@ I Introduction
Playing the piano requires humans to master contact-rich hand movements dictated by the timing and tone they intend to produce. This mastery is not learned quickly but through extensive practice, which requires humans to control their actions based on the haptic and auditory feedback received with each key pressed on the piano. In addition, human hands are an extraordinary research subject due to their unmatched dexterity, precision, and adaptability. These capabilities have evolved to perform various tasks, from delicate manipulations like sewing to powerful tasks like shot-put. This makes human-like hands a natural inspiration for robotic systems. In robotics, much attention has been paid to tasks such as in-hand object rotation, where the goal is to develop and optimize object manipulation through intricate movements [1, 14, 7, 8]. These studies often achieve efficient manipulation by exploiting unnatural and sometimes non-anthropomorphic motions to improve control and precision. Although such techniques yield impressive results, they often diverge from the natural movements of the human hand, which could limit their applicability to tasks that require more human-like actions.
@@
Figure 1: In this work, we demonstrate a proof-of-concept for learning to play piano with a real world robot. To achieve this, we employed a multi-finger robot hand and a Sim2Real approach. Experimental results show that the robot can learn to play several simple pieces successfully, after training exclusively in simulation.
Playing a musical instrument like the piano is an exceptional example of human-like hand function. Piano playing demands a high level of motor control, coordination, and precision, all while closely aligning with the natural movements of human hands. This makes it an ideal scenario for exploring a Simulation to Real World (Sim2Real) transfer, where the objective is to train an agent in simulation capable of performing in the real world. Our work focuses on the challenges of this transfer. We provide interesting insights into the nature of playing piano and actively encourage researchers to participate in this field of study. We believe that by mastering piano playing in the real world, we can bridge the gap between the abstract, often unnatural solutions of robotic manipulation and the fluid, natural motions required to mirror the human approaches to complex tasks more closely.
In this work, we train a reinforcement learning (RL) policy in simulation that is applicable in the real world. The agent’s task is to play simple songs, such as the children’s song ”Twinkle Twinkle Little Star”, with a multi-finger Allegro hand on a real world piano. Our work is based on the open-source training environment RoboPianist-Suite [17]. However, since our primary objective is to deploy the learned policy in the real world, we made numerous changes at the system level. Moreover, due to the changes in the environment, the reward function originally used in the RoboPianist environment was no longer applicable. This pushed us to design a new reward function that is applicable to hands with an arbitrary number of fingers and does not depend on pre-known fingering annotations. To bridge the Sim2Real gap, we utilized domain randomization (DR) of simulation parameters, such as joint dampening, friction coefficients, and spring stiffness.
This work contributes in multiple ways:
  • By providing a proof-of-concept, we want to encourage the community to adopt piano playing as a compelling benchmark towards human-level manipulation.
  • By analyzing three separate execution modes, we gain a deep understanding of the challenges for playing piano in the real world. These observations can be used to develop more capable models.
  • Our model showcases the success with common RL techniques. The results can be used as a baseline to evaluate alternative approaches.
  • Our codebase is open-sourced, which provides researchers a headstart for implementing and verifying new approaches for playing piano in the real world.
II Related work
Understanding the role of robotic systems in playing musical instruments goes a long way back to when highly sophisticated robots such as Takanishi’s Anthropomorphic Flutist Robot [11] that could mimic human respiratory systems were utilized to play the flute. Similarly, early works concerning robots playing the piano indicate the extensive effort made by researchers who built humanoid-like robots that leveraged precise finger-arm movement for rendering tunes [3]. Further works also explored the development of intelligent algorithms [6] and control systems [5] for anthropomorphic robots designed to play the piano. Although path paving, most of the works described heavily relied upon pre-programmed movements, limiting the robot’s ability to adjust to musical styles or improvise on existing scores. Therefore, we move away from traditional methods to a more learnable and robust methodology. Instead, we train the policy in the simulation to play a musical piece without any hard-coded or pre-programmed knowledge, including fingering annotations.
Moreover, playing the piano is a challenging benchmark for dexterity and the contributions in this field are very limited. The contribution of [10] focused on using inverse kinematics to play piano with Shadow hands in the real world, which does not utilize machine learning. In contrast, the contribution by Qian et al. [9], used imitation learning based on YouTube videos to play piano in a simulated environment. Research with robots and piano is novel, and initial works in simulation have demonstrated that learned agents can play simple songs and chords in a simulated environment [16]. Those RL approaches in simulation elaborate on utilizing touch feedback from vision-based tactile sensors [4, 15], to teach a robot hand to play the piano [16]. With the introduction of RoboPianist [17], it was shown that by leveraging the RL algorithms, an agent can be trained to play advanced music with two simulated Shadow hands. Further, the RoboPianist environment was used to build a dataset containing $\approx$1 million trajectories of bimanual piano playing in simulation [18], which can be used to train an imitation learning model. Unlike most prior works, which focus on developing a paradigm to teach robot hands to play in simulation, we focus on translating this learned knowledge into a real world environment.
Sim2Real in robotics has seen widespread research, with a primary focus directed to inducing natural and stable finger gaits and human-like dexterity onto the robots [8]. In our work, we aim to replicate similar human-like learnability during the simulation phase, where we utilize domain randomization to help reinforce the knowledge of the robot and, at the same time, provide scope for generalizing on different musical scores in the real world.
III Learning to Play Piano in the Real World
Hardware Configuration. The hardware setup is shown in figure 2(b). We use a UFACTORY xArm7 to move the end effector along a specified axis parallel to the piano. This mimics the movement of the human forearm. Attached to the end effector is a left-sided Allegro hand v4.0. It consists of 4 fingers with 4 joints per finger, resulting in a total of 16 joints. The thumb joints are not used in this contribution because they can barely reach the piano, leading to 12 active joints. Each of the fingertips is replaced with a 3D-printed finger of approximately the dimensions of a human finger. This is necessary because the default fingertips of the Allegro hand are wider than a key, which makes it impossible to play only one note at a time. Our piano is an M-Audio Keystation 49e MIDI Keyboard, which contains four octaves. Every keypress and every release triggers a MIDI event, which is forwarded to the host computer.
@@ III Learning to Play Piano in the Real World
Figure 2: Comparison of the simulated training environment to the real world. The real world environment consists of a multi-finger Wonik’s Allegro hand mounted on a Ufactory xArm7 robot arm and an M-Audio Keystation 49e MIDI keyboard. The fingertips of the Allegro hand are replaced with thinner 3D-printed fingertips that fit the dimensions of the piano to allow pressing a single key at a time.
Simulation Environment. The environment is based on the RoboPianist environment [17], which was built using the Mujoco physics engine [12] with dm_control [13] as a wrapper for RL tasks. The scene of the simulation is displayed in Fig. 2(a). It consists of the Allegro hand with 3D-printed fingertips and a model of the M-Audio Keystation piano. The hand is modeled analogues to the real world with 12 active joints, 4 per finger. To move the hand parallel to the piano, we introduced another sliding joint for horizontal movement. The piano itself consists of 49 keys and 4 octaves.
@@
Used joints of the Allegro Hand rad 1212
@@
Pressed keys one-hot-encoded 4949
Intended keys one-hot-encoded 4949
Intended keys of future timesteps one-hot-encoded 549=2455*49=245
Total observation space mixed $\boldsymbol{356}$
@@
Markov-Decision-Process. The action space consists of 13 continuous actions. One action per joint of the hand and one additional action for the wrist movement, parallel to the piano. The observation space is 356-dimensional and is shown in Table I. Note that the observations of the keys are discrete. This makes playing the piano more demanding because the agent cannot hover with the fingers on the keys to observe the environment, as described in the RoboPianist [17]. Unfortunately, it is necessary because the MIDI standard used by the piano does not include partially pressed keys. Another contrast to the RoboPianist environment [17] is that the observation space does not include any fingering annotations. This is due to the Allegro hand having only 4 fingers, which makes regular ”human” fingering not applicable. The missing fingering annotations also has a negative impact on the exploration. For those reasons, a carefully chosen reward function is even more crucial to succeed in the exploration.

### v3

@@ Abstract
Towards the grand challenge of achieving human-level manipulation in robots, playing piano is a compelling testbed that requires strategic, precise, and flowing movements. Over the years, several works demonstrated hand-designed controllers on real world piano playing, while other works evaluated robot learning approaches on simulated piano playing. In this work, we develop the first piano playing robotic system that makes use of learning approaches while also being deployed on a real world dexterous robot. Specifically, we use a Sim2Real2Sim approach where we iteratively alternate between training policies in simulation, deploying the policies in the real world, and use the collected real world data to update the parameters of the simulator. Using this approach we demonstrate that the robot can learn to play several piano pieces (including Are You Sleeping, Happy Birthday, Ode To Joy, and Twinkle Twinkle Little Star) in the real world accurately, reaching an average F1-score of $0.881$. By providing this proof-of-concept, we want to encourage the community to adopt piano playing as a compelling benchmark towards human-level manipulation in the real world. We open-source our code and show additional videos at www.lasr.org/research/learning-to-play-piano.
@@ I Introduction
Playing the piano requires humans to master contact-rich hand movements dictated by the timing and tone they intend to produce. This mastery is not learned quickly but through extensive practice, which requires humans to control their hands based on the haptic and auditory feedback received with each key pressed on the piano. In addition, human hands are an extraordinary research subject due to their unmatched dexterity, precision, and adaptability. These capabilities have evolved to perform various tasks, from delicate manipulations like sewing to powerful tasks like shot-put. This makes human-like hands a natural inspiration for robotic systems. In robotics, much attention has been paid to tasks such as in-hand object rotation, where the goal is to develop and optimize object manipulation through intricate movements [Handa2024, wang2024pens, openai2019, Qi2022]. These studies often achieve efficient manipulation by exploiting unnatural and sometimes non-anthropomorphic motions to improve control and precision. Although such techniques yield impressive results, they often diverge from the natural movements of the human hand, which could limit their applicability to tasks that require more human-like motions. This stands in strong contrast to the task of playing a musical instrument like the piano. While piano playing demands a high level of motor control, coordination, and precision it also requires close alignment with the natural movements of human hands.
@@
Figure 1: In this work, we demonstrate a proof-of-concept for learning to play piano with a real world robot. To achieve this, we employed a multi-finger robot hand and a Sim2Real2Sim approach. Experimental results show that the robot can learn to play several simple pieces successfully, after training exclusively in simulation.
Multiple works already explored RL agents for playing piano in simulation [Xu2022, Zakka2023, qian2024, Zhao2024, chen2025, huang2025]. However, we are the first to deploy a learned piano playing agent on a physical robot. Our contribution focuses on transferring the learned strategies in simulation into the real world. This is a demanding task due to the inaccuracies within the training environments and the real world (i.e., Sim2Real gap). This Sim2Real gap can be reduced by tuning the simulation parameters to mimic the real world as close as possible and by training a robust policy. Our approach uses an iterative Sim2Real2Sim approach that utilizes observations of the real world to optimize the training environment. Additionally we use domain randomization during the training to make our learned policy more robust. This enables us to play recognizable songs (without any modification of the tablatures), such as Are You Sleeping, Happy Birthday, Ode To Joy, and Twinkle Twinkle Little Star, with a multi-finger Allegro hand successfully on a real world piano. By providing this proof-of-concept, we want to encourage the community to adopt piano playing as a compelling benchmark towards human-level manipulation.
@@
Figure 2: Comparison of the simulated training environment to the real world setup. The real world environment consists of a multi-finger Allegro hand v4.0 mounted on an Ufactory xArm7 robot arm and an M-Audio Keystation 49e MIDI keyboard. The fingertips of the Allegro hand are replaced with smaller 3D-printed fingertips to allow pressing a single key at a time.
Overall, our contributions are:
  • We introduce the first learning-based system capable of real world piano playing.
  • Our codebase is open-sourced and additional videos are provided on our website. This reduces the initial costs of research in this area and allows researchers to rapidly verify new approaches.
  • We demonstrate that a Sim2Real2Sim is a promising path to explore for future piano playing contributions.
  • Our results can be used as a baseline to evaluate alternative approaches.
  • We provide several important insights into the nature and challenges of playing piano in the real world.
II Related work
Musical Robots. Understanding the role of robotic systems in playing musical instruments goes a long way back to when highly sophisticated robots such as Takanishi’s Anthropomorphic Flutist Robot [takanishi1996] were utilized to play the flute. Similarly, early works concerning robots playing the piano indicate the extensive effort made by researchers who built humanoid-like robots that leveraged precise finger-arm movement for rendering tunes [Kato1987]. More recent works explored the application of robotic hands that are either specifically designed to play the piano [li2013, jen-chang2014] or utilize general purpose hands [scholz2019]. Although path paving, these works heavily relied upon pre-programmed movements, limiting the robot’s ability to adjust to musical styles or improvise on existing scores. This stands in contrast to the contributions utilizing machine learning algorithms to play piano.
Learned Piano Playing. Playing the piano is a challenging benchmark for dexterity that requires precise and flowing movements. The first RL based approach of playing piano was published in 2022 [Xu2022]. The contribution showed that it is possible to play simple notes with correct velocity in simulation by utilizing vision based tactile sensors. With the introduction of RoboPianist [Zakka2023] in 2023, it was shown that a RL-agent can be trained to play advanced bi-manual music. However, the policy needed to be trained for each song individually and manual fingering annotations are required. Those limitations motivate the contribution of Zhao et al. [Zhao2024] in 2024. They proposed fingering as an optimal transport problem [gabriel2019, villani2006] enabling the agent to efficiently discover fingering during training and removing the need for manual finger annotations while also being applicable to other morphologies. To further improve the multi-song performance multiple contributions utilized imitation learning successfully [Zhao2024, qian2024, chen2025]. However, unlike prior works, which focused entirely on developing paradigms to play in simulation, we focused on translating the learned knowledge into the real world.
Sim2Real Transfer. The goal of a Sim2Real transfer is to deploy a policy trained in the simulation in the real world. Unfortunately, the naive approach without Sim2Real transfer techniques leads to a substantial drop in performance. This drop in performance is referred to as the Sim2Real gap. Over time there were multiple methods proposed to minimize the Sim2Real gap e.g. residual physics learning (RPL) [sontakke2023, zeng2020], rapid motor adaptation (RMA) [Qi2022], domain randomization (DR) [Handa2024, tobin2017, wang2024pens] and Sim2Real2Sim [chang2020, lim2022]. They function by either reducing the differences between the simulator and the real world (Sim2Real2Sim, RPL) or by making the model more robust to variations of the environment (DR, RMA). Our work merges both approaches by integrating Sim2Real2Sim and DR.
III Learning to Play Piano in the Real World
@@
Used joints of the Allegro hand rad $12$
@@
Previous joint positions mixed $12+1=13$
Currently pressed keys one-hot-encoded $49$
Music sheet one-hot-encoded 649=2946*49=294
Total observation space $\boldsymbol{369}$
@@
Hardware Configuration. The hardware setup we used is shown in Fig. 2(b). Our setup consists of an Ufactory xArm7 to move the end effector parallel to the piano, replicating the movement of the human forearm. Attached to the end effector is an Allegro Hand v4.0. It consists of 4 fingers with 4 joints per finger. The thumb joints are not used in this contribution because they can not reliably reach the piano, leading to 12 active joints. The fingertips are replaced with 3D-printed fingers of approximately human dimensions, since the default fingertips are wider than a piano key and would prevent single-key presses. Our piano is an M-Audio Keystation 49e MIDI Keyboard, which contains four octaves. Every keypress and every release triggers a MIDI event, which is forwarded to the host computer to allow for observation of the currently pressed keys.
Figure 3: Rendering of the piano keys in our physics simulator. The yellow boundaries are fixed in place and simulate a wall, to prevent the agent from sliding to an adjacent key without lifting the finger. This is discussed in more detail in Section IV-A.
Simulation Environment. Our training environment was built upon the RoboPianist-Suite [Zakka2023], which is based on the Mujoco physics engine [Todorov2012]. The scene of the simulation is displayed in Fig. 2(a). It consists of the Allegro hand with custom fingertips and a model of the piano. Furthermore, we added additional boundaries between the white keys (see Fig. 3) to achieve an environment that mimics the real world more closely. These boundaries are discussed in more detail in Section IV-A. All joints of the simulation use a PD controller with maximum jerk a low pass filter and damping. This enables us to fine-tune the control parameters of all actuators individually during our Sim2Real2Sim workflow.
@@
Markov-Decision-Process. The action space consists of 13 continuous actions. One action per joint of the hand and one additional action controlling the forearm movement. The 369-dimensional observation space is shown in Table I. Note that the observations of the keys are discrete. This makes playing the piano more demanding because the agent cannot hover with the fingers on the keys to verify the correct finger positions, as described in the RoboPianist [Zakka2023]. Unfortunately, it is necessary because the MIDI standard used by the piano does not include partially pressed keys. Further, the observation space does not include any manual fingering annotations. Instead we encourage the agent to use efficient fingering by adding a component to the reward formulation.
Reward Design. We will briefly discuss each component of our reward function $r_{\text{total}}=r_{\text{energy}}+r_{\Delta\text{-action}}+r_{\text{keypress}}+r_{\text{optimal\_transport}}+r_{\text{collision}}$ in this section. The $r_{\text{energy}}$ reward penalizes fast and forceful movements. It avoids oscillations that don’t lead to higher rewards and reduces the applied force to reduce the risk of damaging the real world setup. The $r_{\Delta\text{-action}}$ is a penalty proportional to the difference between consecutive actions to achieve smooth control signals. The $r_{\text{keypress}}$ reward encourages the agent to press the correct keys. To improve the fingering of our robot we added $r_{\text{optimal\_transport}}$, which is adopted from [chen2025]. However, instead of using the cartesian distance as costs we use the one dimensional difference in the horizontal axis to obtain the correct fingertip. Finally, the reward is calculated with respect to the cartesian distance. The motivation for this design is to avoid reusing the same finger just because it has a lower height compared to another finger that is higher up but already above the correct key. Lastly, we added $r_{\text{collision}}$ which is $0.5$ if there is no collision or only collisions between the piano and fingertips, otherwise it is 0. This leads to faster convergence as discussed in Section IV-B.
Evaluation Metrics. To quantify the success of the agent, we use the F1-score $F_{1}=2*(\text{recall}^{-1}+\text{precision}^{-1})^{-1}$, where recall addresses the question how reliable the agent presses the correct keys and precision addresses the question how reliable the agent avoids pressing wrong keys [Zakka2023].
Figure 4: Diagram of our Sim2Real2Sim pipeline. Real world observations from an initial action sequence are used to optimize the simulator. Then a piano-playing policy is trained in simulation and deployed on the robot to generate new data for further simulation refinement. The process is explained in more detail in Section 4.
Sim2Real2Sim. Our workflow is visualized in Fig. 4. We start with an initial action sequence consisting of hand opening and closing motions. This sequence is executed on the real robot (step 1) to collect observations used to optimize the simulation environment using Bayesian Optimization (step 2). The updated environment is then used to train a piano-playing policy (step 3), which is executed on the real robot (step 4) to generate new observations and actions, enabling iterative refinement of the simulation (returning to step 2).
Domain Randomization. To further decrease the Sim2Real gap, we utilize domain randomization of the height of the piano, the starting position of the hand, the dampening of the joints, the applied force of each actuator, the threshold of a key to be considered pressed, the stiffness of the springs holding the keys and the friction between the keys and the fingertips.
Model Training. Our agent consists of three hidden layers with 128 nodes each. The training uses DroQ [Hiraoka2022], an off-policy soft actor-critic algorithm. A new policy needs to be trained for each song, which takes about $3\text{\,}\mathrm{h}$ $20\text{\,}\mathrm{min}$ for $3e6$ epochs.
IV Experimental Results
Figure 5: We analyze the effect of fixed boundaries, referred to as fences, between keys on the training convergence and on the Sim2Real gap. The environment with fences learns slightly faster, but ultimately both environments converge to similar performance. However, the real world evaluation shows an improved F1-score and therefore a smaller Sim2Real gap for the policy trained with fences compared to the fence-free environment. This leads us to the conclusion that adding fences is beneficial, since it reduces the differences between the training environment and the real world.
In the following, we present a series of experiments designed to analyze key design choices of our approach and to evaluate its real world applicability. First, we investigate whether adding physical constraints in simulation, in the form of fences between piano keys, improves training efficiency and Sim2Real transfer. Second, we study the influence of a fingertip collision penalty in the reward formulation on the training and the resulting control strategy. Finally, we assess the real world performance and generalization capability of our approach across multiple songs by comparing the performance achieved in simulation and in the real world.
IV-A Is it beneficial to use fences between the white keys? (Fig. 5)
Setup. In this experiment we compare the training progression and the Sim2Real gap of two simulation environments. The first environment includes narrow fixed boundaries between the white keys, referred to as ”fences” (see Fig. 3). The second environment does not have fences. We compare the convergence during training and the achieved F1-score in the real world.
Results. During training the environment with fences progressed slightly faster compared to the environment without fences. Ultimately both environments converged to an F1-score of $0.946$. In the real world the model without fences achieves an F1-score of $0.840$, leading to a Sim2Real gap of $11.2\%$. In contrast, the model with fences achieved an F1-score of $0.881$, leading to an almost halved Sim2Real gap of $6.8\%$.
Discussion. The results indicate that, in our setup, it is beneficial for the training and real world performance to add fences between the keys. A possible explanation for the faster progression during training is that adding fences makes it simpler to press one key per finger and to hold it pressed without triggering the neighboring keys. Our understanding of the reduced Sim2Real gap is, that the environment with fences makes the simulation behave more similar to the real world. In the real world, applying lateral force to the fingertip while pressing a key does not activate adjacent keys. However, in our fence-free simulation, soft-body physics can cause unintended neighboring key presses. Adding fences forces the agent to completely lift its finger during transitions, better reflecting real world behavior and reducing the risk of getting stuck between keys.
IV-B How does a fingertip collision penalty affect the training? (Fig. 6)
Setup. In this experiment we compare the reward formulation introduced in Section III with a variation that does not include $r_{\text{collision}}$. We observe the progression of the F1-score during training.
Results. The F1-score of the models trained with the collision penalty show a rapid initial performance increase before converging. In contrast, the F1-score of the models without the penalty increases linearly at a slower rate. The gap between both reward formulations is closed at about $\approx 2e6$ episodes and both rewards ultimately converge to the same value.
Discussion. The component $r_{\text{collision}}$ is maximized by playing exclusively with the fingertips. Both formulations converging to this strategy independently indicates that $r_{\text{collision}}$ does not have an influence on the learned strategy. However, the models without the collision penalty require significant more time learning to use the fingertips properly instead of forcing the fingertips inside of the piano by exploiting the soft body physics. This enables us to conclude that the collision penalty primarily accelerates learning rather than altering the final strategy. By discouraging unrealistic collisions early in training, it guides the policy towards a physically plausible fingertip usage. Once this behavior is learned, the reward term becomes constant and no longer influences the optimal policy, explaining the similar performance.
IV-C How does our approach perform in the real world? (Fig. 7)
Setup. To evaluate the generalizability of our approach across songs we tested 5 simple songs: Are You Sleeping, Happy Birthday, Ode To Joy, Twinkle Twinkle Little Star and the C-Major Scale. For each song we obtain the performance achieved in simulation and in the real world to calculate the Sim2Real gap.
Results. All songs reached an F1-score between $0.937$ and $0.964$ in simulation and between $0.821$ and $0.919$ in the real world, however the Sim2Real gap differs significantly from $2\%$ up to $12.3\%$. The average F1-score reached in simulation is $0.946$ and in the real world $0.881$. Videos showcasing the performance of each song are displayed on our website www.lasr.org/research/learning-to-play-piano.
Discussion. The results show that the agent was able to play all 5 songs successfully. The differences in the Sim2Real gap are based on the characteristics of each song e.g. Happy Birthday containing a black key and a faster note sequence in Twinkle Twinkle Little Star. This also indicates that, although our method generalizes to simple songs, more advanced songs continue to pose challenges for the agent.
Figure 6: The diagram compares the training progression of the F1-score with and without the fingertip collision penalty $r_{\text{collision}}$ in the reward formulation. Including the penalty leads to a rapid initial performance increase, while omitting it results in slower, near-linear improvement. Both reward formulations eventually converge to similar final performance.
Figure 7: This experiment is an evaluation of the Sim2Real performance across all 5 songs. All songs are played successfully with an F1-score of $\geq 0.821$ in the real world. However, the Sim2Real gap differs significantly due to the inherent challenges of each song.
V Limitations and Future Work
In this work, we share our findings on how to teach a robot to play simple songs on a real piano. We demonstrated a proof-of-concept that utilized Sim2Real2Sim to play piano in the real world. However, our system possesses several limitations that could be investigated in future research:
Vision Modality. With the current modalities the agent almost entirely relies on its proprioception, which makes perceiving the real world hard. The agent needs to learn the position of each key during training and blindly trust into this knowledge while playing in the real world. To perceive an inaccurate alignment of the fingertip, the agent needs to make a wrong keypress. An additional vision modality would allow the robot to correct itself before pressing the key, potentially improving the overall accuracy.
Tactile Sensors. Another limitation of our work is the lack of additional sensing modalities, including touch sensing. Prior works [Xu2022, nakahara2025] indicate that touch can enable the robot to perceive additional information about the environment. We considered integrating existing touch sensing solutions, such as DIGIT [Lambeta2020]. However, these sensors are larger than a human finger, and resulted in pressing two keys at once. We believe that integrating future generations of tactile sensors that are smaller would be a valuable research topic.
Advanced Fingering. Our work focuses on the Allegro hand which does not allow for finger abduction. This significantly limits the keys each finger can reach without moving the whole hand and it prevents the thumb from reaching the piano. Using a robotic hand that allows for abduction would enable to agent to use more advanced fingering. Further, in human piano playing, the wrist takes a key role to extend the range of each finger for certain fingering. In our work the orientation of the wrist is fixed further reducing the range of each finger. These hardware limitations resulted in movement of the whole hand at the level of the arm. Future research could investigate ways to minimize these movements for more natural results.
Advanced Songs. This work focuses on five simple songs. Playing more advanced songs with faster melodies that require rapid movements is the natural continuation in this direction. This also mimics the progression of a young pianist that learns to play piano with simple songs and one hand before focusing on advanced songs.
Bimanual Playing. Many songs require the use of two hands. This also introduces the challenge of coordinating twice the number of joints and enables more complex fingering techniques. However, currently our hardware setup is limited to a single hand. In the future, we hope to extend to the bimanual case and investigate how to learn more complex piano movements, thus being able to increase the overall difficulty of the played songs.
VI Conclusion
In this work, we showcase the first learned approach for playing piano with a real world robotic hand. Specifically, we employ Sim2Real2Sim where we iterate in the real world before training the final deployed policy. To mimic the real world more accurately, we further introduced boundaries between the white keys in simulation. These boundaries can be easily adopted by other researchers to increase the accurately of their simulations to the real world. Our experimental results demonstrate that our model can play several simple pieces successfully in the real world. Nonetheless, there is still a long way before we can achieve robots with the same level of proficiency and elegance as human pianists. Beyond the artistic raison d’être, we believe that playing piano is a relevant and challenging benchmark for dexterous contact-rich manipulation, and advances in this task can be more widely relevant for the whole manipulation community.
Acknowledgments
We thank Ken Nakahara for the support and tips regarding operating the robotics and throughout the project. This work was supported by the Robotics Institute Germany (RIG), by the German Research Foundation (DFG) under the Cluster of Excellence CARE: Climate-Neutral And Resource-Efficient Construction (EXC 3115) project number 533767731, by the Bundesministerium für Bildung und Forschung (BMBF) and German Academic Exchange Service (DAAD) in project 57616814 (SECAI, School of Embedded and Composite AI), by the project ”Genius Robot” (01IS24083) funded by the Federal Ministry of Education and Research (BMBF), .
Appendix
During the development of our system we also experimented with Hybrid Execution as another Sim2Real transfer technique. Although this was path paving for us, we eventually focused on Sim2Real2Sim in this contribution. In the future we consider combining our Sim2Real2Sim approach with Hybrid Execution to further improve our results. In this section we will describe our preliminary results using Hybrid Execution.
Appendix A Our Hybrid Execution Setup
[…]
Simulation Environment Differences. The simulation environment does not use fences (see Fig. 3). Further, instead of using a pd controllers, we use mujoco position actuators with joint damping to actuate the joints. Besides those differences the environment is analogous.
Markov-Decision-Process. The action space still consists of 13 continuous actions. One action per joint of the hand and one additional action for the wrist movement, parallel to the piano. The observation space however is only 356-dimensional, since it does not include the previous joint positions (compare table I).